OpenStudy (anonymous):

Derivative of F(x) = 35 cos [ 2pi/365 (x-220) ] + 60? I have no idea where to begin...

OpenStudy (anonymous):

lol. ffaaaauuu.....

OpenStudy (anonymous):

I don't know if this is too easy, but the derivative of cos(x) = -sin(x)

OpenStudy (anonymous):

Suppose \[ h(k) = a \cos ( b(x-220)) + c\\ h'(x) =- a b \sin( b(x-220)) \\ \] What are a, b and c in your problem?

OpenStudy (anonymous):

A = 35 B = 2pi/365 C = 220

OpenStudy (anonymous):

Replace and you are done

OpenStudy (anonymous):

Look at C again.

OpenStudy (anonymous):

Oh, that's so simple! Thanks so much!! :)

OpenStudy (anonymous):

c? umm is it supposed to be -220?

OpenStudy (anonymous):

C=60

OpenStudy (anonymous):

it's the constant that goes away with the derivative

OpenStudy (anonymous):

really? isn't 60 something different? like D?

OpenStudy (anonymous):

220 I did not replace with a constant

OpenStudy (anonymous):

No it is not

OpenStudy (anonymous):

Like jbell0385 said, isn't 60 a constant and become 0 when you find its derivative?

OpenStudy (anonymous):

Yeah, I think that's what eliassaab did.

OpenStudy (anonymous):

Your answer should look like \[ F'(x)=-\frac{14}{73} \pi \sin \left(\frac{2}{365} \pi (x-220)\right) \]

OpenStudy (anonymous):

Yes, it looks like that except i didn't simplify 35*(2pi/365). Maybe i should :)

OpenStudy (anonymous):

Yes simplify

OpenStudy (anonymous):

Okay :) Thanks you guys!

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

Solved me a big headache :)